3. A commercial Airline flies frequently between Los Angeles and Chicago. One of the jets flew at and average speed of 640 mph from Los Angeles to Chicago. On the return trip due to turbulance the jet averaged only 512 mph, and the return trip took 48 minutes longer.
a) model the mathematical relationship.
b) How many hours was the flight from Chicago to Los Angeles?
c) What is the distance between Chicago and Los Angeles?

Answer by ankor@dixie-net.com(22740)

3. A commercial Airline flies frequently between Los Angeles and Chicago. One of the jets flew at and average speed of 640 mph from Los Angeles to Chicago. On the return trip due to turbulance the jet averaged only 512 mph, and the return trip took 48 minutes longer.

a) model the mathematical relationship.
Dist = speed * time.
Convert 48 mi hrs: 48/60 = .8 hrs; let t = time for the LA/Chi trip
Distance LA/Chi = Distance Chi/LA
640t = 512(t+.8)
:
b) How many hours was the flight from Chicago to Los Angeles?
640t = 512(t+.8)
640t = 512t + 409.6
640t - 512t = 409.6
128t = 409.6
t = 409.6/128
t = 3.2 hrs, time LA/Chi
Chi/LA = 3.2 + .8 = 4 hrs
:
c) What is the distance between Chicago and Los Angeles?
Dist = speed * time
Dist = 512 * 4
Dist = 2048 miles
:
Check our distance solution using the LA/Chi trip
Dist = 640 * 3.2
Dist = 2048 miles also